Calculate the 99% confidence interval for the following sample.

Sample: 7.9, 8.3, 8.4, 9.6, 7.7, 8.1, 6.8, 7.5, 8.6, 8, 7.8, 7.4, 8.4, 8.9, 8.5, 9.4, 6.9, 7.7.

Assume normality of the data.

Enter the left confidence interval border (round to 3 decimal places):  

Enter the right confidence interval border (round to 3 decimal places):
Answer #1

Assuming normality the 99% confidence interval is given by

Where

Upper Limit : 8.620219

Lower Limit : 7.590892

R code : (This can also run in an online compiler)

x=c(7.9, 8.3, 8.4, 9.6, 7.7, 8.1, 6.8, 7.5, 8.6, 8, 7.8, 7.4, 8.4, 8.9, 8.5, 9.4, 6.9, 7.7)
n=length(x)
s=sd(x)
m=mean(x)
ul= m + s/ sqrt(n) * qt(0.995,n-1)
ll= m - s/ sqrt(n) * qt(0.995,n-1)
ul
ll